Prev Next

SW&IT Meeting Minutes 2019-07-15

Attendees: Andrew, Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every, Do not edit this table, use entries in smartsheets.

News from management

  • kitchen/lounge cleaning rota. details to be discussed.
  • Discussions about the state of documentation in general, and possible ways forward.
  • zoom 3-5 additions welcomed
  • IBM acquires RH for $34B

Previous Action Items

Open topics

Team reports

MB

  • OC, WPC, Management
  • sciarc meeting
  • planning activities
  • CH induction
next
  • vacation

DE

Next
  • Ketiv transfer stuff
  • Script work
  • Vault Office client?

YZ

  • Still working on the LBC skyflats wiki program updates.
  • Getting back to AO-Cravat project (trying to find a time to briefly meet with Al Conrad if possible).
  • LBC/MODS issues:
    • IT#7764: "MODS "updatepointing" command". Requested information was provided and next the OSU team should be able to proceed (maybe reassigning it to Rick).
    • IT#7774: "MODS preset timeout". Rick investigated and added new notes. They know the causes, a comprehensive fix is difficult, so right now maybe just a simple fix to increase the timeout value.
    • IT#7469: "mods1data high disk usage". Discussed with Olga in detail. We are planning to investigate more on how the archival process works, as well as on determining whether an automated process for MODS image deletion is required.
    • Will merge sciops-lbcfpia and sciops-lbcwrs on github into one repo (since that is how the current software is organized and used on obs machine).
    • Checking to make sure the sciops-lbttools repo on github is up-to-date. Some SciOps people may still have local changes uncommitted. Will discuss with them to see if they need to merge back to the main repository.
    • Helping Olga with some of her MODS plotting utilities.
  • Continue to gather remote room/obs machine information.
  • Did some safty training courses.

IS

Last week
  • OT
    • Created a set of scripts that parses the otscripting http log files, extracting scripts and metadata. Then a metadata csv file gets generated. Uncover some issues with MODS scripts not populating the PROPID key and possibly generating dup proposal ids, when scripts are generated in the mountain. Will work on a fix for that.
    • #IT7441 subwindowing for LBC chips
    • Kept working on the OT TPE submenus: fixed toggling offset display between a "per node" basis or all.
This week:
  • OT
    • Finish OT TPE submenu work
    • Look at the script prop id issue, possibly fix it.
    • #IT7074, #IT7192 Issues related to ARGOS and ARGOS guide stars and also the way OT handles errors.

XZ

Last week
  • Discuss "jumping" actutor results with Guido
  • ASMs summer shutdown
  • Given the repair list for this summer's DX ASM wor
  • Build SOUL simulator
Next week:
  • Work on the DX ASM on the summit

PK

  • new engineer starting -- order parts, etc.
  • start zoom for Petr.
Last week
  • IT7743 and IT7695 follow-ups - cleaned git tree of Ice 3.7 changes, plan is to merge Ice 3.7 once ready. All work done locally and pushed to github/pkubanek/tcs; I plan to force push changes to github/LBTO/tcs soon
    • removed Ice 3.7 commits from master and other branches (git rebase -i, d(rop) Ice changes)
    • make sure all changes are in IT7695 branch
    • rebase WIP braches (IT7753??,..) to sit on top of the master branch
    • renamed IT7695 branch to Ice3.7
  • safety training
  • TCS presentation for CIIRC (Czech Institute of Informatics, Robotics and Cybernetics), will share it once done (mostly off work time)
  • vacation

this week:
  • finish Ice3.7 separation
  • CIIRC presentation (Wednesday) (off work time)
  • pointing (TPK) presentation for Al Conrad's summer school (off work time, Monday)

AC

  • IT 6405/7705 - tested wda script to make more atomic. Tried basic file manipulation but symptoms still exist
  • FACSUM testing
next:
  • continue wda and FACSUM build/testing
  • update AzCam MagExposure script and commit to repo
  • solicit AGw telemetry requirements

SH

  • SPOI obs, colo, blah.
  • iLocater networking help
  • RPMs
  • Given possible causes for the actuators
  • Archive firewall open with UITS
  • Clean up telemetry on mountain (slow -- IT 7753).
  • Request for TCS3 removal
next:
  • Google LDAP -- how can I use it?
  • remove tcs3
  • de-crappify my room

CH

  • Monnit telemetry
    • POC: Peter Gray
    • Goal: liberate the sensor tm from imonnit.com via REST API.
    • Monnit's PHP example dumps JSON to console.
    • Will expand upon example for weekly cron download, perhaps in python.
    • Thereafter store as hdf5. I've taken lib-telemetry for a spin for the h5 example.
  • Alarm Handling System (AHS, aka CHAOS?)
    • POC: Matthieu Bec, Petr Kubanek
    • Goal: overhaul/consolidate disparate alerting tools.
    • Will pull existing X11 and web GUI to simulate locally.
    • Will pull copies of known non-AHS alerting scripts.
    • Will poll other PIs to seek out unknown alerting scripts.
Topic revision: r20 - 05 Aug 2019, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback